# Through the Lens of Human-Human Collaboration: A Configurable Research Platform for Exploring Human-Agent Collaboration

### v1

@@ Abstract.
Intelligent systems have traditionally been designed as tools rather than collaborators, often lacking critical characteristics that collaboration partnerships require. Recent advances in large language model (LLM) agents open new opportunities for human-LLM-agent collaboration by enabling natural communication and various social and cognitive behaviors. Yet it remains unclear whether principles of computer-mediated collaboration established in HCI and CSCW persist, change, or fail when humans collaborate with LLM agents. To support systematic investigations of these questions, we introduce an open and configurable research platform for HCI researchers [^1] Our work’s homepage is available at https://human-agent-collab.hailab.io.. The platform’s modular design allows seamless adaptation of classic CSCW experiments and manipulation of theory-grounded interaction controls. We demonstrate the platform’s effectiveness and usability through two case studies: (1) re-implementing the classic human-human-collaboration task Shape Factory as a between-subject human-agent-collaboration experiment with 16 participants, and (2) a participatory cognitive walkthrough with five HCI researchers to refine workflows and interfaces for experiment setup and analysis.
human-agent collaboration, LLM agent, remote collaboration, research platform, open-science
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Human-centered computing Interactive systems and toolsccs: Human-centered computing HCI design and evaluation methods
Figure 1. The core architecture of our research platform, which encompasses four key modules: researcher interface allows researchers to configure experiment setup, monitor live sessions, and perform result analysis; participant interface enables participants to engage in the experiment with other participants and LLM agents; Agent Context Protocol ensures standardized agent integration with the experiment controller; experiment controller manages experiment states and actions updates, and synchronize all the interaction logs into the database.
@@ 1. Introduction
A persistent gap highlighted by the HCI and CSCW community is that intelligent systems are predominantly perceived as “tools,” rather than mutual collaborators with humans (Grudin, 2018; Shneiderman, 2022; Suchman, 1987). Despite their capacity to support humans in data analysis (Wang et al., 2019; Guo et al., 2024), design work (Oh et al., 2018; Lee et al., 2022), and decision-makings (Cai et al., 2019; Zhang et al., 2024), these systems typically lack the mutual awareness (Dourish and Bellotti, 1992; Gutwin and Greenberg, 2002), adaptivity (Allen et al., 1999; Amershi et al., 2019), accountability (Liao et al., 2020; Hoffman et al., 2023), and interdependence (Seeber et al., 2020; Zhang et al., 2021) that characterize effective partnerships. This deficiency often prevents users of intelligent systems to develop trust (Jian et al., 2000; Hoff and Bashir, 2015; Vereschak et al., 2021; Hancock et al., 2011), shared understanding (Van Rensburg et al., 2022; Clark and Brennan, 1991; Clark, 1996), or effective coordination (Schmidt and Bannon, 1992; Olson and Olson, 2000; Hinds and Bailey, 2003), which can further impede the quality of outcomes (Ashktorab et al., 2020; Zheng et al., 2023b), diminish human agency (Tapal et al., 2017; Kim et al., 2024), and undermine reliance in high-stakes situations (Parasuraman and Riley, 1997; Lee and See, 2004; Wang et al., 2020; Vaccaro et al., 2024).
@@
The advent of LLMs has shifted the design space for intelligent systems (Brown et al., 2020; Bubeck et al., 2023). LLMs demonstrate the capacity to engage in fluid communication with humans (Sharma et al., 2023; Yang et al., 2024), while advanced prompting techniques further enable human-like reasoning and planning  (Wei et al., 2022; Zhou et al., 2022; Wang et al., 2023a). Research on LLM agents (Park et al., 2023, 2024; Shanahan et al., 2023), which are LLM systems capable of exhibiting complex, human-like behaviors to solve tasks, shows these agents [^2] In this paper, we use ”LLM agent” and ”agent” interchangeably (Chen et al., 2025a). can yield distinct, believable cognitive and social behaviors, particularly through role-playing the provided persona profiles. These capabilities promise new paradigms of human-agent collaboration, where humans treat these LLM agents as mutual partners rather than mere tools. Yet the fundamental social and cognitive demands of effective collaboration remain (Seeber et al., 2020; Olson and Olson, 2013). The long-standing goal of treating machines as teammates requires moving beyond performance measurement to examine how social (Nass and Moon, 2000; Zhang et al., 2021), communicative (Liang et al., 2019; Zhang et al., 2023), and cognitive behaviors (Schelble et al., 2022; Ashktorab et al., 2020) co-evolve. To achieve flexible and mutual human-agent collaboration, prior work identified that we need interfaces (Amershi et al., 2019; Liao et al., 2020) and interaction paradigms (Wang et al., 2020; He et al., 2024) that help LLM agents communicate intent (Clark and Brennan, 1991; Liang et al., 2019), negotiate decisions (Straus and McGrath, 1994; Bos et al., 2002), manage shared context (Cramton, 2001; Gutwin and Greenberg, 2002; Kraut et al., 2003), and signal responsibility (Erickson and Kellogg, 2000; Hoffman et al., 2023; Liao et al., 2020) in ways humans can readily perceive and act upon.
@@
To inform the design of interactions in the context of human-agent collaboration, our vision is that LLM agents should resemble remote human collaborators. Both LLM agents and remote human collaborators rely on structured channels such as text or voice, and both lack social cues (e.g., facial expressions, tone) that augment in-person interactions (Short et al., 1976; Walther, 1996; Olson and Olson, 2000). While this analogy is limited by differences in accountability, knowledge grounding, and identity persistence, it motivates designing interfaces and interactions for human–LLM-agent collaboration by drawing on decades of CSCW and HCI research on remote collaboration. Empirical and theoretical research has systematically identified key interaction variables in remote collaboration: communication modalities shape the richness of information exchange (Daft and Lengel, 1986; Straus, 1997; Bos et al., 2002), workspace awareness supports coordination (Dourish and Bellotti, 1992; Gutwin and Greenberg, 2002), and information transparency fosters trust and mutual understanding (Cramton, 2001; Hinds and Bailey, 2003). Recent studies confirmed that these foundational principles remain relevant as new agent capabilities emerge (Zhang et al., 2023; He et al., 2024; Lu et al., 2025). Grounding human-LLM agent interaction in this body of research allows HCI researchers to investigate which principles and challenges persist, which are shifted, and what new issues arise as boundaries between remote human collaborators and LLM agents blur.
@@
To address this gap, we present an open, configurable research platform for conducting reproducible, controlled experiments on human-LLM-agent collaboration. Our platform design draws inspiration from the Shape Factory experiment (Bos et al., 2004), a classic CSCW experiment for analyzing group dynamics between collocated and remote collaborators. The Shape Factory experiment supports rigorous manipulation of communication modality, awareness, social framing, and resource interdependence. Our platform enables HCI researchers to adapt classic experimental paradigms for human-agent collaboration research, such as strategic collaboration tasks (e.g., DayTrader (Bos et al., 2002)), collaborative decision-making tasks (e.g., Essay Ranking (Zheng et al., 2023b)), and collaborative task-solving (e.g., Passcode (Gero et al., 2020). The platform architecture features four key components, as shown in Fig 1:
@@
Critically, the platform allows systematic manipulation of theory-grounded interaction controls to investigate their impact on collaborative dynamics, such as communication modality (Straus, 1997; Bos et al., 2002), awareness dashboards (Gutwin and Greenberg, 2002; Dourish and Bellotti, 1992), and social framing (Bos et al., 2004; Hinds and Bailey, 2003). Researchers can independently manipulate these controls and further customize the functionalities to meet more nuanced needs using the platform’s open-source code.
We demonstrate our platform’s utility through a two-part evaluation. First, we validate the usability and effectiveness of our platform for controlled experiments through two case studies that re-implemented the Shape Factory human-human collaboration experiment (six-person team) into a human-agent collaboration experiment (one person and five LLM agents as a team) (Bos et al., 2004). A total of 16 participants took part in a crossed, between-subjects design with counterbalancing in both case studies, where we manipulated a key interaction control in each. The goal was to confirm that our platform could faithfully re-implement a classic experimental paradigm and produce measurable differences in participant behaviors and outcomes. Analysis of behavioral logs and post-study surveys confirmed the platform’s ability to capture significant differences across conditions; for instance, varying communication modality led to distinct patterns in negotiation frequency and team performance, as well as substantial differences in perceived trust and workspace awareness, which is aligned with established findings in collaborative work. Second, we conducted a participatory cognitive walkthrough with five HCI researchers on human-AI collaboration to evaluate the researcher interface for experiment setup and data analysis. Their feedback informed iterative improvements to the workflow and interface design. Collectively, these evaluations validate our platform as a methodological foundation for the HCI community to build a systematic, evidence-based understanding of human-agent collaboration.
@@ 2. Related Work
We group and discuss relevant literature into four bodies of work: (1) the shift in HCI from AI as a tool to AI as a partner, (2) how LLMs change the collaboration landscape and what still remains unsolved, (3) CSCW research on remote collaboration that offers a lens for human-LLM agent collaboration, and (4) experimental paradigms and platforms for collaboration studies, and why a theory-driven, extensible platform for human–agent experiments is still missing but of significant importance to the HCI and CSCW community.
2.1. Human-AI Collaboration: From Tools to Partners
Intelligent systems have often been designed and perceived as tools, rather than partners that share goals, coordinate actions, and carry responsibility with people (Grudin, 2018; Shneiderman, 2022; Suchman, 1987). Research on human-AI collaboration highlights several gaps that impede intelligent systems from establishing partnerships with humans.
Effective partnerships depend on awareness (i.e., the ability to know who is doing what, where, and when (Dourish and Bellotti, 1992; Gutwin and Greenberg, 2002)) and on common ground (i.e., a shared understanding of context and references (Clark and Brennan, 1991) in collaborative systems). When systems obscure state, history, and intent, collaborators have to guess and repair misunderstandings, raising coordination cost and slows progress (Dourish and Bellotti, 1992; Gutwin and Greenberg, 2002; Clark and Brennan, 1991; Schmidt and Bannon, 1992; Clark, 1996). Besides, adaptivity and initiative are critical pillars to support mixed-initiative in collaboration, where a system should be able to decide when to act, when to wait, and how to make its moves understandable (Allen et al., 1999; Amershi et al., 2019).
Many intelligent systems, particularly before the emergence of LLMs, only offered template-based dialog or single-shot commands and failed to adapt to users’ evolving goals nor proactively offer the right help at the right time (Amershi et al., 2019). In addition, accountability was also underscored to allow human collaborators to ask a system why a recommendation was made, solicit supporting evidence, and examine uncertainty; otherwise, users may experience mis-calibrated trust and reliance (Liao et al., 2020; Hoffman et al., 2023; Jian et al., 2000; Lee and See, 2004; Hancock et al., 2011). Subsequently, trust mis-calibration often will result in poorer joint outcomes and reluctance to use systems in higher-stakes work (Parasuraman and Riley, 1997; Wang et al., 2020; Vaccaro et al., 2024). Moreover, collaborative outcomes highly depend on interdependence, where team members should engage with each other to negotiate roles and divide work (Seeber et al., 2020). Many work, however, treat systems as isolated predictors or recommenders and leave little room for negotiating and clarifying labor distribution, which will weaken shared understanding (Van Rensburg et al., 2022; Clark, 1996) and make coordination inaccessible (Schmidt and Bannon, 1992). When intelligent systems cannot fulfill the need as a partner and are instead recognized as tools, studies reported lower joint quality (Ashktorab et al., 2020; Zheng et al., 2023b) and reduced sense of agency and engagement (Tapal et al., 2017; Kim et al., 2024).
2.2. The New Landscape of Human-AI Collaboration in the Era of Large Language Model-based Agent
@@
The emergence of LLMs sheds light on new opportunities for the landscape of LLM-based intelligent systems as mutual collaborators of humans. In particular, LLMs enable natural language as a shared medium that allows users to state goals, constraints, and references, while LLM systems can ask clarifying questions and propose plans both in free-form language (Brown et al., 2020; Bubeck et al., 2023; Sharma et al., 2023; Yang et al., 2024). This capability shifts interaction beyond rigid command structures into fluid, open-ended conversations. Furthermore, advanced prompting techniques, such as chain-of-thought (Wei et al., 2022; Kojima et al., 2022), enable LLMs to perform step-wise reasoning and planning for their predictions. These techniques allow visible forms of system accountability and initiative that people can read and negotiate.
@@
Building on these technical achievements, researchers have developed “LLM role-playing agents,” which adopt personas and goals to simulate complex social and cognitive behaviors over extended interactions (Park et al., 2023, 2024; Shanahan et al., 2023). Applications are rapidly expanding: LLM agents have been explored as human surrogates for automated UX testings (Lu et al., 2025), and as tools for simulating populations and replicating classic findings in social science (Argyle et al., 2023; Aher et al., 2023).
@@
Despite these advances of LLM agents, the fundamental social and cognitive demands for effective human-agent collaboration remain. The key lies in the design of interfaces (Amershi et al., 2019; Liao et al., 2020) and interaction paradigms (Wang et al., 2020; He et al., 2024) between humans and LLM agents to establish the collaboration needs identified in classic CSCW research. However, HCI researchers currently lack a dedicated research platform that allow systematic mapping of classic collaboration insights onto the interface features of LLM agents. Such a platform is essential for investigating how collaboration dynamics unfold in this new era of human–AI partnerships.
2.3. Human-Human Remote Collaboration as an Analytical Lens of Exploring Human-Agent Collaboration

### v2

@@ Abstract.
Intelligent systems have traditionally been designed as tools rather than collaborators, often lacking critical characteristics that collaboration partnerships require. Recent advances in large language model (LLM) agents open new opportunities for human-LLM-agent collaboration by enabling natural communication and various social and cognitive behaviors. Yet it remains unclear whether principles of computer-mediated collaboration established in HCI and CSCW persist, change, or fail when humans collaborate with LLM agents. To support systematic investigations of these questions, we introduce an open and configurable research platform for HCI researchers [^1] Our platform is available at https://github.com/neuhai/human-agent-collab.. The platform’s modular design allows seamless adaptation of classic CSCW experiments and manipulation of theory-grounded interaction controls. We demonstrate the platform’s research efficacy and usability through three case studies: (1) two Shape Factory experiments for resource negotiation with 16 participants, (2) one Hidden Profile experiment for information pooling with 16 participants, and (3) a participatory cognitive walkthrough with five HCI researchers to refine workflows of researcher interface for experiment setup and analysis.
human-AI collaboration, human-agent collaboration, CSCW, LLM agents, remote collaboration, research platform, controlled experiments, open-science, workspace awareness, common ground
journalyear: 2026copyright: ccconference: Proceedings of the 2026 CHI Conference on Human Factors in Computing Systems; April 13–17, 2026; Barcelona, Spainbooktitle: Proceedings of the 2026 CHI Conference on Human Factors in Computing Systems (CHI ’26), April 13–17, 2026, Barcelona, Spaindoi: 10.1145/3772318.3790879isbn: 979-8-4007-2278-3/2026/04ccs: Human-centered computing Interactive systems and toolsccs: Human-centered computing HCI design and evaluation methods
Figure 1. The core architecture of our research platform, which encompasses four key modules: researcher interface allows researchers to configure experiment setup, monitor live sessions, and analyze result; participant interface enables human participants to engage with other participants and LLM agents; Agent Context Protocol ensures standardized agent integration; experiment controller manages experiment states and actions updates, and synchronize all the interaction logs into the database.
Fig. 1. The figure shows the core architecture of a research platform with four modules. A researcher interface (top left) supports experiment setup, live monitoring, and result analysis. A participant interface (top right) lets humans and LLM agents interact in the experiment (e.g., chat/trade, status, and dashboard). An experiment controller (center) manages experiment state and participant actions, and synchronizes all interaction logs to a database (bottom left). An Agent Context Protocol (ACP) (bottom right) standardizes communication between LLM agents and the experiment controller.
@@ 1. Introduction
The HCI and CSCW communities have long highlighted a gap where intelligent systems are predominantly perceived as “tools,” rather than mutual collaborators with humans (Grudin, 2018; Shneiderman, 2022; Suchman, 1987). Despite their capacity to support humans in data analysis (Wang et al., 2019; Guo et al., 2024), design work (Oh et al., 2018; Lee et al., 2022), and decision-makings (Cai et al., 2019; Zhang et al., 2024b), these systems typically lack the mutual awareness (Dourish and Bellotti, 1992; Gutwin and Greenberg, 2002), adaptivity (Allen et al., 1999; Amershi et al., 2019), accountability (Liao et al., 2020; Hoffman et al., 2023), and interdependence (Seeber et al., 2020; Zhang et al., 2021) that characterize effective partnerships. This deficiency often impedes trust (Jian et al., 2000; Hoff and Bashir, 2015; Vereschak et al., 2021; Hancock et al., 2011), shared understanding (Van Rensburg et al., 2022; Clark and Brennan, 1991; Clark, 1996), or effective coordination (Schmidt and Bannon, 1992; Olson and Olson, 2000; Hinds and Bailey, 2003), which can further limits the quality of outcomes (Ashktorab et al., 2020; Zheng et al., 2023a), diminish human agency (Tapal et al., 2017; Kim et al., 2024), and undermine reliance in high-stakes situations (Parasuraman and Riley, 1997; Lee and See, 2004; Wang et al., 2020; Vaccaro et al., 2024).
@@
The advent of LLMs has shifted the design space for intelligent systems (Brown et al., 2020; Bubeck et al., 2023). LLMs demonstrate the capacity to engage in fluid communication with humans (Sharma et al., 2023; Yang et al., 2024), while advanced prompting techniques further enable human-like reasoning and planning  (Wei et al., 2022; Zhou et al., 2022; Wang et al., 2023). Research on LLM agents (Park et al., 2023, 2024; Shanahan et al., 2023), which are LLM systems capable of exhibiting complex, human-like behaviors to solve tasks, shows these agents [^2] In this paper, we use ”LLM agent” and ”agent” interchangeably (Chen et al., 2025a). can yield distinct, believable cognitive and social behaviors, particularly through role-playing the provided persona profiles. These capabilities promise new paradigms of human-agent collaboration, where humans treat these LLM agents as mutual partners rather than mere tools.
However, deploying these agents into real-world teams introduces critical socio-technical challenges. In high-stakes domains such as healthcare triage or crisis management, failures to establish common ground or to signal uncertainty can degrade collective performance and erode accountability (Cai et al., 2019; Zhang et al., 2024b). For instance, does an ”assertive” agent facilitate faster decision-making, or does it suppress critical thinking and information sharing among collaborators? Addressing these design problems requires isolating interaction variables like communication modality and agent responsiveness in human-agent collaboration. Yet, investigating these dynamics requires moving beyond simple performance metrics to examine how social (Nass and Moon, 2000; Zhang et al., 2021), communicative (Liang et al., 2019; Zhang et al., 2023), and cognitive behaviors (Schelble et al., 2022; Ashktorab et al., 2020) co-evolve in controlled settings.
@@
To inform the design of interactions in the context of human-agent collaboration, our vision is that LLM agents should resemble remote human collaborators. Both LLM agents and remote human collaborators rely on structured channels such as text or voice, and both lack social cues (e.g., facial expressions, tone) that augment in-person interactions (Short et al., 1976; Walther, 1996; Olson and Olson, 2000). While this analogy is limited by differences in accountability, knowledge grounding, and identity persistence, it motivates designing interfaces and interactions for human–LLM-agent collaboration by drawing on decades of CSCW and HCI research on remote collaboration. Empirical and theoretical research has systematically identified key interaction variables in remote collaboration: communication modalities shape the richness of information exchange (Daft and Lengel, 1986; Straus, 1997; Bos et al., 2002), workspace awareness supports coordination (Dourish and Bellotti, 1992; Gutwin and Greenberg, 2002), and information transparency fosters trust and mutual understanding (Cramton, 2001; Hinds and Bailey, 2003). Recent studies confirmed that these foundational principles remain relevant as new agent capabilities emerge (Zhang et al., 2023; He et al., 2024; Lu et al., 2025). Grounding human-LLM agent interaction in this body of research enables HCI researchers to investigate which principles and challenges persist, which shift, and what new issues arise as the boundaries between remote human collaborators and LLM agents blur.
@@
To address this gap, we present an open, configurable research platform for conducting reproducible, controlled experiments on human-LLM-agent collaboration. Our platform design draws inspiration from the Shape Factory experiment (Bos et al., 2004), a classic CSCW experiment for analyzing group dynamics between collocated and remote collaborators. The Shape Factory experiment supports rigorous manipulation of communication modality, awareness, social framing, and resource interdependence. Our platform enables HCI researchers to adapt classic experimental paradigms for human-agent collaboration research, such as strategic collaboration tasks (e.g., DayTrader (Bos et al., 2002)), collaborative decision-making tasks (e.g., Hidden Profile (Stasser and Titus, 1985, 1987) and Essay Ranking (Zheng et al., 2023a)), and collaborative task-solving (e.g., Passcode (Gero et al., 2020)). The platform architecture features four key components, as shown in Fig 1:
@@
Critically, the platform allows systematic manipulation of theory-grounded interaction controls to investigate their impact on collaborative dynamics, such as communication modality (Straus, 1997; Bos et al., 2002), awareness dashboards (Gutwin and Greenberg, 2002; Dourish and Bellotti, 1992), and social framing (Bos et al., 2004; Hinds and Bailey, 2003). Researchers can independently manipulate and customize these controls to meet nuanced needs using the platform’s open-source code.
We demonstrate our platform’s utility through a two-part evaluation. First, we validate the research efficacy of our platform for controlled experiments through multi-phase case studies. We implemented the Shape Factory experiment (Bos et al., 2004) to investigate how communication channels (chat v.s. no chat) and workspace awareness (with v.s. without dashboard) influence resource negotiation in a mixed human-agent team. This phase employs a crossed, between-subjects design with 16 participants. Subsequently, we implemented the Hidden Profile experiment (Stasser and Titus, 1985, 1987) to examine how agent responsiveness (passive v.s. proactive) affects information pooling and collaborative decision-making quality. This phase employs a between-subjects design with 16 participants. These studies confirmed that the platform could faithfully re-implement classic paradigms and capture significant divergence in individual and collaborative behaviors, collaboration outcome, and perception caused by different interaction designs. Second, we conducted a participatory cognitive walkthrough with five HCI researchers on human-AI collaboration to evaluate the researcher interface for experiment setup and data analysis. Their feedback informed iterative improvements to workflow and interface design. Collectively, these evaluations validate our platform as a methodological foundation for the HCI community to advance a systematic, evidence-based understanding of human-agent collaboration.
@@ 2. Related Work
We review four lines of work that motivate the need for a theory-driven, extensible platform for human–agent collaboration studies.
2.1. Barriers to Treating Human-AI Collaboration as a Partnership
Although intelligent systems are increasingly positioned as collaborators rather than mere tools, prior HCI and CSCW research shows several persistent barriers prevent them from functioning as true partners to humans. These barriers revolve around limited awareness and common ground, insufficient mixed-initiative support, and the lack of accountability and interdependence structures needed for coordinated joint work.
Awareness and Common Ground. Effective collaboration requires knowing who is doing what, when, and why (Dourish and Bellotti, 1992; Gutwin and Greenberg, 2002) and maintaining shared references and assumptions (Clark and Brennan, 1991). When systems obscure their state, history, or intent, users must infer or repair breakdowns on their own (Schmidt and Bannon, 1992; Clark, 1996), which could increase coordination costs and slow work progress.
Initiative and Adaptivity. Mixed-initiative collaboration depends on a system’s ability to determine when to act, when to wait, and how to make its moves interpretable (Allen et al., 1999; Amershi et al., 2019). Pre-LLM systems were largely template-driven and unable to adapt to evolving goals or proactively provide the right help (Amershi et al., 2019), making them poor partners in dynamic tasks.
Accountability and Trust Calibration. Partners must be able to explain their opinions and reasoning, justify recommendations, and communicate uncertainty (Liao et al., 2020; Hoffman et al., 2023). Without such accountability, users develop mis-calibrated trust (Jian et al., 2000; Lee and See, 2004; Hancock et al., 2011), which leads to degraded collective outcomes and reluctance to rely on systems in high-stakes tasks (Parasuraman and Riley, 1997; Wang et al., 2020; Vaccaro et al., 2024).
Interdependence and Role Negotiation. Collaboration also depends on interdependence-partners coordinating roles, dividing labor, and negotiating contributions (Seeber et al., 2020). Many intelligent systems, however, operate as isolated predictors or recommenders (Van Rensburg et al., 2022), offering little opportunity for role clarification or shared control (Clark, 1996). As a result, users experience weakened shared understanding and diminished agency and engagement (Ashktorab et al., 2020; Zheng et al., 2023a; Tapal et al., 2017; Kim et al., 2024).
Fig. 2. This figure shows a comparison of existing research platforms for collaboration studies. Top row, from left to right: Share, GroupKit, LabintheWild, and Empirica. Bottom row: oTree, nodeGame, Overcooked, and Pairit. These platforms represent diverse approaches to studying collaboration.
2.2. The New Landscape of Human-AI Collaboration in the Era of LLM Agent
@@
The emergence of LLMs sheds light on new opportunities for the landscape of LLM-based intelligent systems as mutual collaborators of humans. In particular, LLMs enable natural language as a shared medium that allows users to state goals, constraints, and references, while LLM systems can ask clarifying questions and propose plans in free-form language (Brown et al., 2020; Bubeck et al., 2023; Sharma et al., 2023; Yang et al., 2024). This capability shifts interaction beyond rigid command structures into fluid, open-ended conversations. Furthermore, advanced prompting techniques, such as chain-of-thought (Wei et al., 2022; Kojima et al., 2022), enable LLMs to perform step-wise reasoning and planning for their predictions. These techniques allow visible forms of system accountability and initiative that people can read and negotiate.
@@
Building on these technical achievements, researchers have developed “LLM role-playing agents,” which adopt personas and goals to simulate complex social and cognitive behaviors over extended interactions (Park et al., 2023, 2024; Shanahan et al., 2023). Applications are rapidly expanding: LLM agents have been explored as human surrogates for automated UX testing (Lu et al., 2025), and as tools for simulating populations and replicating classic findings in social science (Argyle et al., 2023; Aher et al., 2023).
@@
Despite these advances of LLM agents, the fundamental social and cognitive demands for effective human-agent collaboration remain. The key lies in the design of interfaces (Amershi et al., 2019; Liao et al., 2020) and interaction paradigms (Wang et al., 2020; He et al., 2024) between humans and LLM agents to establish the collaboration needs identified in classic CSCW research. However, HCI researchers currently lack a dedicated research platform that allows systematic mapping of classic collaboration insights onto the interface features of LLM agents. Such a platform is essential for investigating how collaboration dynamics unfold in this new era of human–AI partnerships.
2.3. Remote Human Collaboration as an Analytical Lens of Exploring Human-Agent Collaboration
LLM agents resemble remote human collaborators in important ways: both operate through structured, text-based channels and lack the non-verbal cues that facilitate grounding and intent inference in face-to-face interaction (Short et al., 1976; Walther, 1996; Olson and Olson, 2000). CSCW research on distributed teamwork, therefore, offers a valuable analytical lens for understanding human-agent collaboration challenges.